given a right triangle, use the pythagorean theorem to find a when c=85 and b = 53
i tried this a^2 + b^2 = c^2
a^2+53^2 = 85^2
a^2 + 2809=7225
-2809=-2809
a^2=4416

Found 2 solutions by stanbon, MORUPHOSUOLALE:
Answer by stanbon(75887)

given a right triangle, use the pythagorean theorem to find a when c=85 and b = 53
i tried this a^2 + b^2 = c^2
a^2+53^2 = 85^2
a^2 + 2809=7225
-2809=-2809
a^2=4416
a = 66.45..
Cheers,
Stan H.

Answer by MORUPHOSUOLALE(24)

Hello there,
We know that this is a phytagorean theorem.
I will assume that the lenght of the hypotenus side is 85 since it is true and constant that the hypotenus side is usually the longest of the three sides of any right angle triengle.
Therefore,
a^2=b^2+c^2
85^2=53^2+c^2
7225=2809+c^2
7225-2809=c^2 (by collecting like terms)
4418=c^2
66.5 = c 1.e the sqr of4418.
I hope this helps
Moruph Osu.